Find the locus Β foot of perpendicular from focus to the tangent y^2=4ax
Β 

Grade:12th pass

2 Answers

Vikas TU
7 years ago
Β 
  • The foot of the perpendicular from the focus on any tangent to a parabola lies on the tangent at the vertex.

Equation of the perpendicular to the tangent ty = x + at2 … (1)

From the focus (a, 0) is tx + y = at. Β  Β  Β  Β  Β  Β  Β  Β  Β  Β  Β  Β  Β  Β  Β  … (2)

By adding (1) and (2) we get x = 0.(Since (1 + t2)Β β‰  0)

Hence, the point of intersection of (1) and (2) lies on x = Β 0 i.e. on y-axis. This is also the tangent at the vertex of the parabola.

  • The tangent at any point P on aΒ parabolaΒ bisects the angle between the focal chord through P and the perpendicular from P on the directrix.

?The tangent at P (at2, 2at) is ty = x + at2.

It meets the x-axis at T(–at2, 0).

Hence, from the figure given aboveΒ ST = SA + AT = a (1 + t2).Β 

Also, SP = √(a2(1 + t2)2 + 4a2 t2 ) = a(1 + t2) = ST, so that 

∠MPT = ∠PTS = ∠SPT β‡’ TP bisects ∠SPM.

  • The portion of a tangent to a parabola cut off between the directrix and the curve subtends a right angle at the focus.

Let P(at2, 2a), be a point on the parabola y2Β = 4ax.

Then the equation of tangent at P is ty = x + at2.

Point of intersection of the tangent with the directrix x + a = 0 is (–a, at – a/t).

Now, slope of SP is (2at-0)/(at2-a) = 2t/(t2-1)Β and slope of SK is (at-a/t-0)/(-a-a) = -(t2-1)/2tΒ 

β‡’ (Slope of the SP).(Slope of SK) = –1.

Hence SP is perpendicular to SK i.e. ∠KSP = 90°. 

  • Tangents at the extremities of any focal chord intersect at right angles on the directrix.Β 

?Let P(at2, 2at) and P(at12, 2at1) be the end points of a focal part on the parabola. Then t.t1Β = –1. Equations of the tangents at the point P and the point P’ are ty = x + at2Β and t1y = x + at12Β respectively.

Let these tangents intersects at a point (h, k). Then h = att1Β and k = a(t + t1).

Since the tangents are perpendicular, tt1Β = – 1 β‡’ h – a.

Hence the locus of the point (h, k) is x = –a which is the equation of the directrix.Β 

Riya Chandra
2 years ago
Tangent to parabola is, y = mx +a/m ....(i)
A line perpendicular to tangent and passing from focus (a, 0) is, y =-x/m +a/m ......(ii)
Solving both lines (i) and (ii) β‡’ x = 0Β